OpenStudy (anonymous):

Can someone help me with this question? The topic is Systems of Equations. It says: Bo and Erica are yoga instructors. Between the two of them, they teach 42 yoga classes each week. If Erica teaches 12 fewer than twice as many as Bo, how many classes does each instructor teach per week? 22 Bo; 20 Erica 20 Bo; 22 Erica 15 Bo; 27 Erica 18 Bo; 24 Erica

OpenStudy (lyrae):

b = number of students Bo teaches. e = number of student Erica teaches. Between the two of them, they teach 42 yoga classes each week. b + e = 42 Erica teaches 12 fewer than twice as many as Bo 2b - 12 = e -> 2b - e = 12 You now have the following system to solve b + e = 42 2b - e = 12

OpenStudy (anonymous):

@Lyrae Thanks so much!

OpenStudy (lyrae):

Yw :) And a small correction: b and e are the number of classes they teach, not the amount of students. My mind were somwhere else :P